# Panoramic Out-of-Distribution Segmentation for Autonomous Driving

### v2

@@ V-A Datasets
We train the model on Cityscapes [15] and extra inlier dataset: Mapillary Vistas [83] for outlier exposure fine-tuning. For PanOoS evaluation, we use DenseOoS and QuadOoS benchmarks. Additionally, we also use Road Anomaly [51] to evaluate pinhole-OoS.
  • Cityscapes consists of $2,975$ training and $500$ validation images and contains $19$ categories which are considered as inliers in out-of-distribution segmentation benchmarks, such as road, sky, building, and car etc.
  • Mapillary Vistas is a large-scale street-level image dataset containing $25,000$ high-resolution images with a variety of weather, season, time of day, camera, and viewpoint, providing rich visual information.
  • Road Anomaly is a collection of $60$ web images with diverse anomaly objects on or near the road.
@@ V-B Implementation Details
Closed-Set Training. We follow [29] and adopt the same training recipe and losses without any special design. The difference is that we adopt the ConvNeXt-Large CLIP [75] from OpenCLIP [84], pre-trained on the LAION-2B dataset [85], and freeze its weights. In reference to previous research works [28, 27], only one decoder layer with $100$ queries is used in the transformer decoder. Both dimensions $C_{m}$ and $C_{q}$ are set to $256$. The model is trained for $90k$ iterations using a batch size of $16$ on $4$ NVIDIA GeForce RTX 3090 GPUs. The training is optimized with the AdamW optimizer and weight decay $0.05$. We employ the initial learning rate $1{\times}10^{-4}$ and a polynomial schedule.
Outlier Exposure Fine-tuning. After the model is trained on the closed-set setting, we fine-tune the pixel decoder, $3$-layer MLP, and Linear layers for $30k$ iterations on the inlier set of Cityscapes [15] and Mapillary Vistas [83] using the setting of the closed-set training; AdamW optimizer with $0.05$ weight decay and $1{\times}10^{-4}$ initial learning with polynomial schedule. For every inlier image used in fine-tuning, an object from the MS-COCO samples [69] is uniformly chosen as an outlier and pasted on the image using AnomalyMix [32] with a probability $p_{out}$ (set to $0.3$), which is independent for each image.
@@
PEBAL [32] 16.39 7.59 46.29 11.34 39.25 13.87 23.42
DenseHybrid [33] 24.44 7.74 45.19 19.09 33.86 21.77 20.80
EAM [26] 17.97 6.46 61.58 33.05 79.66 25.51 43.06
Mask2Anomaly [28] 36.28 8.79 48.10 25.93 92.55 31.10 50.67
RbA [27] 51.31 21.87 57.54 34.12 90.34 42.71 56.11
@@
TABLE II: Panoramic out-of-distribution segmentation quantitative evaluation: POS shows significant improvement over baseline per-pixel and mask-transformer-based pinhole-OoS methods on PanOoS. Higher values for AuPRC, and mIoU are better, whereas for FPR95 lower values are better. The best and second best results are bold and underlined, respectively, and the line in-between the table divides the per-pixel architecture and mask-transformer-based methods.
@@
DAFormer [34] 44.53
Trans4PASS (Tiny) [12] 44.93
Trans4PASS (Small) [12] 49.35
Trans4PASS+ (Tiny) [2] 49.29
Trans4PASS+ (Small) [2] 50.11
360SFUDA++ [48] 50.46
@@
V-C Evaluation Metrics
In the context of the Panoramic Out-of-distribution Segmentation (PanOoS) benchmark, we employ three metrics, namely Area under PRC Curve (AuPRC), False Positive Rate at a true positive rate of $95\%$ (FPR95), and mean Intersection-over-Union (mIoU), to evaluate the performance of models. We provide a detailed explanation. For pixel-wise evaluation, $\mathcal{G}{\in}\{\mathcal{G}_{in},\mathcal{G}_{out}\}$ is the annotated ground truth labels for an image containing outliers, where $\mathcal{G}_{in}$ and $\mathcal{G}_{out}$ represents the inlier and outlier labels, respectively. Let $\mathcal{G}_{p}$ denote the model prediction obtained by a threshold $\tau$. The True Positive Rate (TPR), precision, and recall can be written as:
The AuPRC can be approximated as:
$AuPRC\!$ $=\!\int_{\tau}\!precision(\tau)\,recall(\tau).$ (27)
where $\tau_{95}$ is a threshold when TPR is $95\%$.
$mIoU\!=\frac{1}{K}\!\sum_{k=1}^{K}\!\frac{p_{k}\!\cap\!g_{k}}{p_{k}\!\cup\!g_{k}}.$ (29)
V-D Results of Panoramic Out-of-distribution Segmentation
Figure 13: Visual comparisons of PanOoS. We observe that when existing pinhole-OoS methods are transferred to PanOoS, current state-of-the-art methods generally struggle with misidentifying complex background as outliers. Although EAM [26] has a low false positive rate, it incorrectly identifies bicycles as outliers and nearly overlooks the actual outliers. In contrast, Mask2Anomaly [28] and RbA [27] suffer from large false positives in the background region, whereas POS delivers more precise PanOoS across the $360^{\circ}$. Zoom in for a better view.
We deliver PanOoS results on DenseOoS and QuadOoS benchmarks. Two representative types of existing state-of-the-art pinhole-OoS methods, pixel-wise architecture, and mask-transformer-based methods, are benchmarked. As shown in Table II, mask-transformer-based OoS methods demonstrate strong potential. Compared to the best-performing method RbA [27], POS outstrips it by large margins of $34.25\%$ in AuPRC, $21.42\%$ in FPR95, and $5.76\%$ in mIoU on DenseOoS. Furthermore, state-of-the-art panoramic semantic segmentation models [34, 12, 2, 48] are compared (see Table III). Existing panoramic semantic segmentation methods cannot distinguish between in- and out-of-distribution regions and are hindered by outliers, resulting in degraded closed-set segmentation performance. However, POS achieves excellent out-of-distribution segmentation while maintaining outstanding in-distribution segmentation performance, demonstrating the strong potential of prompt distribution learning in omnidirectional scene perception. We further explore PanOoS in unconstrained, unstructured, and dense-motion real-world scenarios, where all baseline methods struggle to make accurate predictions under motion blur, resulting in significant false positive issues. Specifically, pixel-wise methods are not very sensitive to blur but struggle to effectively recognize outliers within the scene, while mask-transformer-based methods tend to produce false positives. POS achieves a good balance, with mean AuPRC and FPR95 improvements of approximately $24\%$ and $13\%$, respectively, on the two benchmarks.
@@ V-D Results of Panoramic Out-of-distribution Segmentation
PEBAL [32] 62.37 28.29
DenseHybrid [33] - -
EAM [26] 69.40 7.70
Mask2Anomaly [28] 79.70 13.45
RbA [27] 85.42 6.92
@@
In Fig. 13, we visualize the PanOoS results produced by POS and other mask-transformer-based methods. Notably, POS effectively suppresses false positives in background regions, especially at the boundaries separating inliers, while better preserving the smoothness of the outlier map compared to other methods, despite sharing the same mask classification training paradigm. Benefiting from the generalization ability of vision-language models, Prompt-based Restoration Attention (PRA) and Bilevel Prompt Distribution Learning (BPDL), POS performs excellently on the two benchmarks for PanOoS, perfectly adapting to the background clutter and pixel distortions caused by the broader FoV of panoramic images.
V-E Results of Pinhole Out-of-distribution Segmentation
To further investigate the segmentation capacity of the proposed POS in the pinhole domain, we evaluate its OoS performance on one standard pinhole-OoS datasets [51]. During the outlier exposure fine-tuning, we do not apply panoramic-oriented disentanglement and fine-tune only the MLP and the linear layers. The results in Table IV indicate that POS excels on Road Anomaly, and achieves precise segmentation of outliers with minimal false positives.
V-F Ablation Study
@@
Each Component in POS. Table V presents the component-wise ablation results of the technical novelties included in POS. Using the image encoder alone as the backbone, the model demonstrates strong potential for omnidirectional visual perception. However, it still struggles with semantic ambiguity and background interference in complex panoramic scenes. By incorporating Prompt-based Restoration Attention (PRA), the model significantly improves pixel-level semantic associations through prompt distribution learning, enhancing panoramic perception. This results in an increase in AuPRC to $83.94\%$ ($7.06\%$) and an improvement in mIoU to $62.50\%$ ($0.77\%$), highlighting the effectiveness in mitigating dense pixel semantic blurring. Further incorporation of Bilevel Prompt Distribution Learning (BPDL) leads to a synergistic optimization: PRA enhances inter-class discrimination through semantic decoupling, while BPDL preserves topological consistency via manifold learning. This combined mechanism enables the model to achieve groundbreaking performance in the PanOoS task: $85.56\%$ AuPRC, $0.45\%$ FPR95, and $63.30\%$ mIoU. Not only does it overcome the coupling challenges of semantic ambiguity and geometric distortion that traditional pinhole-OoS methods cannot deal with, but it also establishes a new benchmark for panoramic scene understanding.
@@
Figure 14: Attention maps of queries in PRA after different query stages, resembling a confusion matrix in which each query has the highest semantic similarity with itself. When querying only with image features, the query semantic information is relatively vague. However, after further interaction with the prompt embeddings, the semantics become clearer, and the self-adaptive correction further mitigates the impact of pixel distortion on semantic decoding.
Prompt-base Restoration Attention. To better understand the effect of Prompt-based Restoration Attention (PRA), we perform an ablation analysis of its structure in Table VI. Experimental results show that the simple use of Masked Attention (MA), which only interacts with image features, cannot adapt to the characteristics of panoramic images, resulting in poor performance across almost all metrics. This demonstrates that prompt-based Cross Attention (CA) effectively captures high-level semantic interactions, alleviating the impact of pixel distortions on semantic decoding. Furthermore, to prevent blind confidence during the learning process, the proposed Self-adaptive Correction (SAC) dynamically adjusts the semantic confidence for each query, leading to more accurate semantic segmentation and avoiding excessively high voting scores for out-of-distribution areas. This is visually substantiated in Fig. 14, where the inter-query attention maps after three stages explicitly manifest the effectiveness of PRA for hierarchical semantic refinement.
@@
Bilevel Prompt Distribution Learning. The proposed BPDL belongs to the domain of metric learning. Unlike traditional methods that rely on manually defined or learned centroid embeddings as class prototypes, BPDL uses semantically rich text embeddings as prototypes. This approach allows for a coarse-grained predefinition of distribution regions for each class. Moreover, BPDL utilizes a bilevel learning framework for distribution optimization: the lower-level optimization refines the inter-class distributions for known classes, while the upper-level optimization adjusts the global distribution space. As shown in Table VII, each component of the $\mathcal{L}_{bpdl}$ contributes to spatial distribution optimization, with the $\mathcal{L}_{ind}$ and $\mathcal{L}_{pixel}$ loss terms leading to particularly notable performance gains (AuPRC: $2.95\%/{+}2.48\%$, mIoU: $0.63\%/{+}0.41\%$). Ablation experiments confirm that integrating all loss components results in optimal performance across all evaluation metrics.
As outlined in Eqs. 9, 18, 19, and 24, the BPDL framework involves four key hyperparameters: inter-class margin $s$, distribution margin $d$, loss weight of $\mathcal{L}_{Dis}$ $\alpha$, and balancing coefficient $\lambda_{bdpl}$. The inter-class margin $s$ controls the trade-off between intra-class embedding compactness and inter-class separability, while the distribution margin $d$ governs the geometric structure of the embedding space. The parameter $\alpha$ adjusts the relative contribution of the $\mathcal{L}_{Dis}$ loss, and $\lambda_{bpdl}$ balances the weights between the $\mathcal{L}_{bpdl}$ loss and other loss terms. Through a systematic sensitivity analysis (Fig. 15), we observe a notable convex trend for the $s$. Both mIoU and AuPRC exhibit convex behavior as functions of $s$, while FPR95 shows a concave trend but remains stable within a fluctuation range of $0.03\%$. Similarly, the distribution margin $d$ follows a similar optimization trend, confirming the synergistic effect of the bilevel distribution constraints. The parameter $\alpha$ regulates the distance between the upper-level class prototype embeddings and the distribution embeddings, enhancing the isolation of out-of-distribution embeddings. However, an excessive emphasis on this constraint can reduce closed-set semantic segmentation performance. Additionally, when $\lambda_{bpdl}$ exceeds a certain threshold ($0.01$), the model tends to focus too much on the out-of-distribution detection task, leading to significant degradation in closed-set segmentation performance (with mIoU decreasing by $0.6\%$), while out-of-distribution detection metrics remain relatively stable.
@@
Figure 15: Experiments of different settings of $s$, $d$, $\alpha$ and $\lambda_{bpdl}$.
Num Iter AuPRC FPR95 mIoU
15kk 79.19 0.49 57.49
20kk 78.22 0.57 60.14
25kk 84.04 0.51 62.71
30kk 85.56 0.45 63.30
35kk 82.08 0.56 63.14
((a))
$p_{out}$ AuPRC FPR95 mIoU
0.1 86.30 0.45 62.47
0.2 85.72 0.45 62.73
0.3 85.56 0.45 63.30
0.4 81.28 0.59 62.17
0.5 80.97 0.53 62.36
((b))
Module AuPRC FPR95 mIoU
w/o 55.66 7.44 54.11
Decoders 84.25 0.50 62.38
Pixel Dec. 85.56 0.45 63.30
Trans. Dec. 46.91 5.58 57.89
((c))
TABLE VIII: Ablation study on outlier exposure fine-tuning.
Outlier Exposure Fine-tuning. We further empirically evaluate the impact of fine-tuning parameters during the optimization of the prompt distribution space. As shown in Table LABEL:tab:ablationIter, performance evaluations on the DenseOoS dataset for PanOoS demonstrate that the model reaches optimal performance after approximately $30k$ iterations, after which overfitting begins. For the outlier data exposure control parameter $p_{out}$, experiments reveal that when $p_{out}{>}0.3$, both closed-set performance and out-of-distribution segmentation performance deteriorate. Based on these results, we selected $0.3$ as the optimal value, balancing outlier detection with closed-set performance.

### v3

@@ V-A Datasets
We train the model on Cityscapes [cordts2016cityscapes] and an extra inlier dataset: Mapillary Vistas [neuhold2017mapillary] for outlier exposure fine-tuning. For PanOoS evaluation, we use DenseOoS and QuadOoS benchmarks. Additionally, we use Road Anomaly [lis2019detecting] to evaluate pinhole-OoS.
  • Cityscapes consists of $2,975$ training and $500$ validation images and contains $19$ categories that are considered inliers in out-of-distribution segmentation benchmarks, such as road, sky, building, and car.
  • Mapillary Vistas is a large-scale street-level image dataset containing $25,000$ high-resolution images with a variety of weather, seasons, times of day, cameras, and viewpoints, providing rich visual information.
  • Road Anomaly is a collection of $60$ web images with diverse anomalous objects on or near the road.
@@ V-B Implementation Details
V-B1 Closed-Set Training
We follow [cheng2022masked] and adopt the same training recipe and losses without any special design. The difference is that we adopt the ConvNeXt-Large CLIP [liu2022convnet] from OpenCLIP [ilharco_gabriel_2021_5143773], pre-trained on the LAION-2B dataset [schuhmann2022laionb], and freeze its weights. In reference to previous research works [rai2024mask2anomaly, nayal2023rba], only one decoder layer with $100$ queries is used in the transformer decoder. Both dimensions $C_{m}$ and $C_{q}$ are set to $256$. The model is trained for $90k$ iterations using a batch size of $16$ on $4$ NVIDIA GeForce RTX 3090 GPUs. The training is optimized with the AdamW optimizer and a weight decay $0.05$. We employ the initial learning rate of $1{\times}10^{-4}$ and a polynomial schedule.
V-B2 Outlier Exposure Fine-tuning
After training in the closed-set setting, we fine-tune the pixel decoder, the $3$-layer MLP, and the linear layers for $30k$ iterations on the inlier set of Cityscapes [cordts2016cityscapes] and Mapillary Vistas [neuhold2017mapillary], using the setting of closed-set training; AdamW optimizer with $0.05$ weight decay and $1{\times}10^{-4}$ initial learning rate with polynomial schedule. For each inlier image used in fine-tuning, an object from the MS-COCO samples [lin2014microsoft] is uniformly chosen as an outlier and pasted on the image using AnomalyMix [tian2022pixel] with a per-image probability $p_{out}$ (set to $0.3$).
V-C Results of Panoramic Out-of-distribution Segmentation
We deliver PanOoS results on the DenseOoS and QuadOoS benchmarks. Two representative types of existing state-of-the-art pinhole-OoS methods, pixel-wise architectures, and mask-transformer-based methods, are benchmarked. As shown in Table II, mask-transformer-based OoS methods demonstrate strong potential. Compared to the best-performing method RbA [nayal2023rba], POS outstrips it by $34.25\%$ in AuPRC, $21.42\%$ in FPR95, and $5.76\%$ in mIoU on DenseOoS. Furthermore, state-of-the-art panoramic semantic segmentation models [hoyer2022daformer, zhang2022bending, zhang2024behind, zheng2024360sfuda++] are compared (see Table III). Existing panoramic semantic segmentation methods cannot distinguish between in- and out-of-distribution regions and are hindered by outliers, resulting in degraded closed-set segmentation performance. However, POS achieves excellent out-of-distribution segmentation while maintaining outstanding closed-set segmentation performance, demonstrating the strong potential of prompt distribution learning in omnidirectional scene perception. We further explore PanOoS in unconstrained, unstructured, and dense-motion real-world scenarios, where all baseline methods struggle to make accurate predictions under motion blur, resulting in significant false positive issues. Specifically, pixel-wise methods are not very sensitive to blur but struggle to effectively recognize outliers within the scene, while mask-transformer-based methods tend to produce false positives. POS achieves a good balance, with mean AuPRC and FPR95 improvements of approximately $24\%$ and $13\%$, respectively, on the two benchmarks. Benefiting from the generalization ability of vision-language models, Prompt-based Restoration Attention (PRA) and Bilevel Prompt Distribution Learning (BPDL), POS performs excellently on the two benchmarks for PanOoS, perfectly adapting to the background clutter and pixel distortions caused by the broader FoV of panoramic images.
@@ V-C Results of Panoramic Out-of-distribution Segmentation
PEBAL [tian2022pixel] 16.39 7.59 46.29 11.34 39.25 13.87 23.42
DenseHybrid [grcic2022densehybrid] 24.44 7.74 45.19 19.09 33.86 21.77 20.80
EAM [grcic2023advantages] 17.97 6.46 61.58 33.05 79.66 25.51 43.06
Mask2Anomaly [rai2024mask2anomaly] 36.28 8.79 48.10 25.93 92.55 31.10 50.67
RbA [nayal2023rba] 51.31 21.87 57.54 34.12 90.34 42.71 56.11
@@
TABLE II: Panoramic out-of-distribution segmentation quantitative evaluation. POS shows significant improvement over baseline per-pixel and mask-transformer-based pinhole-OoS methods on PanOoS. Higher values for AuPRC and mIoU are better, whereas for FPR95 lower values are better. The best and second best results are bold and underlined, respectively, and the line in-between the table divides the per-pixel architectures and mask-transformer-based methods.
@@
DAFormer [hoyer2022daformer] 44.53
Trans4PASS (Tiny) [zhang2022bending] 44.93
Trans4PASS (Small) [zhang2022bending] 49.35
Trans4PASS+ (Tiny) [zhang2024behind] 49.29
Trans4PASS+ (Small) [zhang2024behind] 50.11
360SFUDA++ [zheng2024360sfuda++] 50.46
@@
PEBAL [tian2022pixel] 62.37 28.29
DenseHybrid [grcic2022densehybrid] - -
EAM [grcic2023advantages] 69.40 7.70
Mask2Anomaly [rai2024mask2anomaly] 79.70 13.45
RbA [nayal2023rba] 85.42 6.92
@@
V-D Results of Pinhole Out-of-distribution Segmentation
To further investigate the segmentation capacity of the proposed POS in the pinhole domain, we evaluate its OoS performance on one standard pinhole-OoS datasets [lis2019detecting]. During the outlier exposure fine-tuning, we do not apply panoramic-oriented disentanglement and fine-tune only the MLP and the linear layers. The results in Table IV indicate that POS excels on Road Anomaly, and achieves precise segmentation of outliers with minimal false positives.
V-E Ablation Study
@@
V-E1 Each Component in POS
Table V presents the component-wise ablation results of the technical novelties included in POS. Using the image encoder alone as the backbone, the model demonstrates strong potential for omnidirectional visual perception. However, it still struggles with semantic ambiguity and background interference in complex panoramic scenes. By incorporating Prompt-based Restoration Attention (PRA), the model significantly improves pixel-level semantic associations through prompt distribution learning, enhancing panoramic perception. This results in an increase in AuPRC to $83.94\%$ ($7.06\%$) and an improvement in mIoU to $62.50\%$ ($0.77\%$), highlighting the effectiveness in mitigating dense pixel semantic blurring. Further incorporation of Bilevel Prompt Distribution Learning (BPDL) leads to a synergistic optimization: PRA enhances inter-class discrimination through semantic decoupling, while BPDL preserves topological consistency via manifold learning. This combined mechanism enables the model to achieve groundbreaking performance in the PanOoS task: $85.56\%$ AuPRC, $0.45\%$ FPR95, and $63.30\%$ mIoU. Not only does it overcome the coupling challenges of semantic ambiguity and geometric distortion that traditional pinhole-OoS methods cannot deal with, but it also establishes a new benchmark for panoramic scene understanding.
@@ V-E1 Each Component in POS
Figure 11: Attention maps of queries in PRA after different query stages, resembling a confusion matrix in which each query has the highest semantic similarity with itself. When querying only with image features, the semantic information of queries is relatively vague. However, after further interaction with the prompt embeddings, the semantics become clearer, and the self-adaptive correction further mitigates the impact of pixel distortion on semantic decoding.
V-E2 Prompt-base Restoration Attention
To better understand the effect of Prompt-based Restoration Attention (PRA), we perform an ablation analysis of its structure in Table VI. Experimental results show that using Masked Attention (MA) alone, which interacts only with image features, cannot adapt to the characteristics of panoramic images, resulting in poor performance across almost all metrics. This demonstrates that prompt-based Cross Attention (CA) effectively captures high-level semantic interactions, alleviating the impact of pixel distortions on semantic decoding. Furthermore, to prevent blind confidence during the learning process, the proposed Self-adaptive Correction (SAC) dynamically adjusts the semantic confidence for each query, yielding more accurate semantic segmentation and avoiding excessively high voting scores for out-of-distribution areas. This is visually substantiated in Fig. 11, where the inter-query attention maps after three stages explicitly manifest the effectiveness of PRA for hierarchical semantic refinement.
@@ V-E2 Prompt-base Restoration Attention
Figure 12: Sensitivity analysis of $s$, $d$, $\alpha$ and $\lambda_{bpdl}$.
V-E3 Bilevel Prompt Distribution Learning
The proposed BPDL belongs to the domain of metric learning. Unlike traditional methods that rely on manually defined or learned centroid embeddings as class prototypes, BPDL uses semantically rich text embeddings as prototypes. This approach allows for a coarse-grained predefinition of distribution regions for each class. Moreover, BPDL utilizes a bilevel learning framework for distribution optimization: the lower-level optimization refines the inter-class distributions for known classes, while the upper-level optimization adjusts the global distribution space. As shown in Table VII, each component of the $\mathcal{L}_{bpdl}$ contributes to feature distribution optimization, with the $\mathcal{L}_{ind}$ and $\mathcal{L}_{pixel}$ loss terms leading to particularly notable performance gains (AuPRC: $2.95\%/{+}2.48\%$, mIoU: $0.63\%/{+}0.41\%$). Ablation experiments confirm that integrating all loss components results in optimal performance across all evaluation metrics.
As outlined in Eqs. 10, 19, 20, and 25, the BPDL framework involves four key hyperparameters: inter-class margin $s$, distribution margin $d$, loss weight of $\mathcal{L}_{Dis}$ $\alpha$, and balancing coefficient $\lambda_{bpdl}$. The inter-class margin $s$ controls the trade-off between intra-class embedding compactness and inter-class separability, while the distribution margin $d$ governs the geometric structure of the embedding space. The parameter $\alpha$ adjusts the relative contribution of the $\mathcal{L}_{Dis}$ loss, and $\lambda_{bpdl}$ balances the weights between the $\mathcal{L}_{bpdl}$ loss and other loss terms. Through a systematic sensitivity analysis (Fig. 12), we observe a notable convex trend for $s$. Both mIoU and AuPRC exhibit convex behavior as functions of $s$, while FPR95 shows a concave trend but remains stable within a fluctuation range of $0.03\%$. Similarly, the distribution margin $d$ follows a similar optimization trend, confirming the synergistic effect of the bilevel distribution constraints. The parameter $\alpha$ regulates the distance between the upper-level class prototype embeddings and the distribution embeddings, enhancing the isolation of out-of-distribution embeddings. However, an excessive emphasis on this constraint can reduce closed-set semantic segmentation performance. Additionally, when $\lambda_{bpdl}$ exceeds a certain threshold ($0.01$), the model tends to focus too much on the out-of-distribution detection task, leading to significant degradation in closed-set segmentation performance (with mIoU decreasing by $0.6\%$), while out-of-distribution detection metrics remain relatively stable.